Solve the following LPP:  

Minimize Z= 120x1+60x2 


Subject to:  

20x1+30x2>= 900  

40x1+30x2>=1200  

x1,x2>=0


1
Expert's answer
2022-01-31T12:09:42-0500

Solution:

Minimize "Z= 120x_1+60x_2"

Subject to:  

"20x_1+30x_2\ge 900 

\\40x_1+30x_2\ge1200 

\\x_1,x_2\ge0"



A(0,40), B(15,20), C(45,0) are corner points.

At A(0,40) , "Z= 120x_1+60x_2 =120(0)+60(40)=2400" (Minimum)

At B(15,20),"Z= 120x_1+60x_2 =120(15)+60(20)=3000"

At C(45,0),"Z= 120x_1+60x_2 =120(45)+60(0)=5400"

Thus, minimum is 2400 at "x_1=0, x_2=40"